# A route to quantum computing through the theory of quantum graphs

### v1

@@ Abstract
Based on our previous works, and in order to relate them with the theory of quantum graphs and the quantum computing principles, we once again try to introduce some newly developed technical structures just by relying on our toy example, i.e. the coordinate ring of $n\times n$ quantum matrix algebra $M_{q}(n)$, and the associated directed locally finite graphs $\mathcal{G}(\Pi_{n})$, and the Cuntz-Krieger $C^{*}$-graph algebras. Meaningly, we introduce a $(4i-6)$-quantum entangled system by using the Cuntz-Krieger $\mathcal{G}(\Pi_{i})$-families associated to the $64i-$ 6 distinct Hamiltonian paths of $\mathcal{G}(\Pi_{i})$, for $i\in\{2,\cdots,n\}$.
@@ 4.1 A route to quantum computing through the theory of quantum graphs
Postulate 4.8 (State Space Postulate).
@@
A linear combination $\phi=c_{1}|0\rangle+\cdots+c_{p}|p-1\rangle$ of the basis vectors of $\mathbb{C}^{p}$, for $c_{1},\cdots,c_{p}\in\mathbb{C}$ will be called a qubit.
@@
In Definition 4.10, $c_{i}$s, for $i\in\{1,\cdots,p\}$, usually (in the literature) will be called amplitudes, and they satisfy in the condition $|c_{1}|^{2}+\cdots+|c_{p}|^{2}=1$, which is the usual normalization condition, it is not too difficult to see that the normalization condition will simply imply $\||\phi\rangle\|=\sqrt{\langle\phi|\phi\rangle}=1$.
@@
So, let us invite $\mathcal{G}(\Pi_{2})$ back to the scene. Take $m=4$ in quantum channel $\Psi_{\mathcal{G}(\Pi_{2})}$. Here, our quantum system will be a combination of two subsystems $\mathbb{E}_{1}$ and $\mathbb{E}_{2}$ with the correspondent states $\phi_{1}$ and $\phi_{2}$, and since in this case, the Hamiltonian paths are a combination of three edges, hence let $\mathbb{C}^{3}$ be our state space $\phi$. Consider $\phi_{1}=\alpha_{1}|0\rangle+\alpha_{2}|1\rangle+\alpha_{3}|2\rangle$ and $\phi_{2}=\beta_{1}|0\rangle+\beta_{2}|1\rangle+\beta_{3}|2\rangle$ such that we have $\sum_{i=1}^{3}\alpha_{i}^{2}=\sum_{i=1}^{3}\beta_{i}^{2}=1$. Not that the only possible options for $\alpha_{i}$s and $\beta_{i}$s are ±$\pm\frac{1}{\sqrt{3}}$.
@@
Our pre-assumed 2-qubit quantum system associated with $\mathcal{G}(\Pi_{2})$ is entangled!
@@
The proof of this proposition is almost a trivial conclusion of the Definition 4.17, because, in any case, $|\phi\rangle$ will never be equal to the tensor product $|\phi_{1}\rangle\otimes|\phi_{2}\rangle$, because amplitudes in $|\phi_{1}\rangle\otimes|\phi_{2}\rangle$ in their best possible state will be equal to ±$\pm\frac{1}{3}$, while in $|\phi\rangle$ we have ±$\pm\frac{1}{2}$. ∎
@@
[1] A quantum state $|\phi\rangle\in(\mathbb{C}^{m})^{\otimes q}$ is a product state if it can be expressed as a tensor product $|\phi_{1}\rangle\otimes\cdots\otimes|\phi_{q}\rangle$ of $q$ 1-qubit states. Otherwise, it is entangled.
@@
In the case of $\mathcal{G}(\Pi_{n})$, the pre-assumed $(4n-6)$-qubit quantum system is entangled.

### v2

@@ Abstract
Based on our previous works, and in order to relate them with the theory of quantum graphs and the quantum computing principles, we once again try to introduce some newly developed technical structures just by relying on our toy example, i.e. the coordinate ring of $n\times n$ quantum matrix algebra $M_{q}(n)$, and the associated directed locally finite graphs $\mathcal{G}(\Pi_{n})$, and the Cuntz-Krieger $C^{*}$-graph algebras. Meaningly, we introduce a $(4i-6)$-qubit quantum system by using the Cuntz-Krieger $\mathcal{G}(\Pi_{i})$-families associated to the $64i-$ 6 distinct Hamiltonian paths of $\mathcal{G}(\Pi_{i})$, for $i\in\{2,\cdots,n\}$.
@@ 4.1 A route to quantum computing through the theory of quantum graphs
Remark 4.7.
Hence, from lemma 4.6, we know that in $\mathcal{G}(\Pi_{n})$ there are $4n-6=2(2n-3)$ distinct Hamiltonian paths with edge degree $n^{2}-1$. But, later on we will notice that we cannot use this as the dimension of our subsystems, but the dimension is very close to this number!
@@
Postulate 4.9 (State Space Postulate).
@@
A linear combination $\phi=c_{1}|0\rangle+c_{2}|1\rangle$ of the basis vectors of $\mathbb{C}^{2}$, for $c_{1},c_{2}\in\mathbb{C}$ will be called a qubit.
@@
In Definition 4.11, $c_{i}$s, for $i\in\{1,2\}$, usually (in the literature) will be called amplitudes, and they satisfy in the condition $|c_{1}|^{2}+|c_{2}|^{2}=1$, which is the usual normalization condition, it is not too difficult to see that the normalization condition will simply imply $\||\phi\rangle\|=\sqrt{\langle\phi|\phi\rangle}=1$.
@@
From remark 4.7, for $i\in\{2,\cdots,n\}$, in $\mathcal{G}(\Pi_{i})$, we know that the dimension of our subsystems are not equal to $i^{2}-1$, which is the edge dimension of the Hamiltonian paths. Instead, let us take $k=i^{2}-1-((i-1)^{2}+1)=2i-3$. Now consider $(\mathbb{C}^{2})^{\otimes 2i-3}$ to be our subsystems of dimension $2^{4i-6}$ divided to the product of $2\times(2i-3)$ 1-qubit states.
Now, let us invite $\mathcal{G}(\Pi_{2})$ back to the scene. Take $m=4$ in quantum channel $\Psi_{\mathcal{G}(\Pi_{2})}$.
Here, our quantum system will be a combination of two subsystems $\mathbb{E}_{1}$ and $\mathbb{E}_{2}$ with the correspondent states $\phi_{1}$ and $\phi_{2}$, and since in this case, the Hamiltonian paths are a combination of three edges, hence let $(\mathbb{C}^{2})^{\otimes 1}\cong\mathbb{C}^{2}$ be our state space $\phi$. Consider $\phi_{1}=\alpha_{1}|0\rangle+\alpha_{2}|1\rangle$ and $\phi_{2}=\beta_{1}|0\rangle+\beta_{2}|1\rangle$ such that we have $\sum_{i=1}^{2}\alpha_{i}^{2}=\sum_{i=1}^{2}\beta_{i}^{2}=1$. Not that the only possible options for $\alpha_{i}$s and $\beta_{i}$s are ±$\pm\frac{1}{\sqrt{2}}$.
@@
Our pre-assumed 2-qubit, and 6-qubit quantum systems associated with $\mathcal{G}(\Pi_{2})$, and $\mathcal{G}(\Pi_{3})$, respectively, are not entangled!
@@
The proof of this proposition is almost a trivial conclusion of the Definition 4.18, because, in any case, $|\phi\rangle$, and $|\psi\rangle$, respectively for $\mathcal{G}(\Pi_{2})$ and $\mathcal{G}(\Pi_{3})$, will always be equal to the tensor product $|\phi_{1}\rangle\otimes|\phi_{2}\rangle$, because in $\mathcal{G}(\Pi_{2})$ the amplitudes $|\phi_{1}\rangle\otimes|\phi_{2}\rangle$ will be equal to ±$\pm\frac{1}{2}$, just the same as in $|\phi\rangle$, and in $\mathcal{G}(\Pi_{3})$ the amplitudes $|\psi_{1}\rangle\otimes|\psi_{2}\rangle$ will be equal to ±$\pm\frac{1}{2^{3}}$, just the same as in $|\psi\rangle$
@@
[1] A quantum state $|\phi\rangle\in(\mathbb{C}^{2})^{\otimes q}$ is a product state if it can be expressed as a tensor product $|\phi_{1}\rangle\otimes\cdots\otimes|\phi_{q}\rangle$ of $q$ 1-qubit states. Otherwise, it is entangled.
@@
In the case of $\mathcal{G}(\Pi_{n})$, the pre-assumed $(4n-6)$-qubit quantum system is not entangled.
@@
Remark 4.20.
If you have already noted that the quantum state of our quantum systems will be $(\mathbb{C}^{2})^{\otimes 2}=\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, $(\mathbb{C}^{2})^{\otimes 6}=(\mathbb{C}^{2})^{\otimes 3}\otimes(\mathbb{C}^{2})^{\otimes 3}$, $(\mathbb{C}^{2})^{\otimes 10}=(\mathbb{C}^{2})^{\otimes 5}\otimes(\mathbb{C}^{2})^{\otimes 5}$, and so on, with dimensions $4,64,1024$ and so on, in order. And hence we might formulate the following claim.
Claim 4.21.
The quantum system arising from $\mathcal{G}(\Pi_{n})$ agrees and improves (in quantum scale) the ordinary classical systems!